# Multi-head committees enable direct uncertainty prediction for atomistic foundation models

### v1

@@ IV Methods
The custom-built models consisted of two message-passing layers, with 32 channels and a maximum tensor order of $l=1$. The multi-layer perceptron in the output blocks of the final layer had 16 nodes in the hidden layer of each output block and used the default SiLU gate. The radial cut-off was set to $6.0\text{\,}\AA$ , resulting in an effective field of view of $12.0\text{\,}\AA$ . The isolated atom energies were always set to the ones specified in the respective datasets. The compositional differences between the committee types lead to different sizes of the full training dataset used for each model. To keep the total number of optimisation steps consistent between the different types of models, we adapted the training parameters in our setups. This also means that each member of the naive committee received as many optimisation steps as the MHC. When investigating different training set sizes, we again ensured that the number of optimisation steps remains constant for all training sets. A weighted mean squared error with a force-to-energy weight ratio of 100 to 1 was used for model optimisation. For the final 25% of the training, the Stochastic Weight Averaging approach with default settings was used to further optimise the energy predictions. The 3BPA models were trained for 5000 optimisation steps, the water models for 2200 steps, and the rMD17 models for 16,000 steps. The output blocks of the adapted foundation model were trained with the same basic parameters for a total of 800,000 steps. To condense the MPtrj dataset, we used an iterative QbC workflow [Schran2020/10.1063/5.0016004]. In each iteration, we sorted the structures based on the maximum disagreement of their force components and added the 100 highest ranking structures to the training set. Afterward, the output heads were retrained for a reduced number of optimisation steps.
@@
Data will be made public on Zenodo when published, DOI xx.xxxx
@@
Acknowledgements.
H.B. and P.S. acknowledge support from the Charles University Grant Agency, project number 248923, and the International Max Planck Research School for Quantum Dynamics and Control. L.L.S. would like to acknowledge support from the UKRI Critical Mass grant, project reference EP/V062654/1, the Isaac Newton Trust, award number G122390 and Wolfson College, Cambridge. O.M. acknowledges support from the Czech Science Foundation, project No. 21-27987S. C.S. acknowledges financial support from the Royal Society, grant number RGS/R2/242614.
Author contributions
Competing Interests
The authors declare no competing interests.
@@ Acknowledgements.
When investigating how well committee disagreement functions as an uncertainty estimation for the true error in the main article, we focused on a graphical evidence and plotted the uncertainty estimate against the actual error. However, there are alternatives to assess uncertainty in a more quantitative way, which we show here. The first is the Pearson correlation coefficient:
$r(\epsilon,\sigma)=\frac{\sum_{A}\left(\epsilon(A)-\overline{\epsilon}\right)\left(\sigma(A)-\overline{\sigma}\right)}{\sqrt{\sum_{A}\left(\epsilon(A)-\overline{\epsilon}\right)^{2}}\sqrt{\sum_{A}\left(\sigma(A)-\overline{\sigma}\right)^{2}}},$ (S1)
which is a measure of correlation between two datasets, in our instance, the error $\epsilon$ and the uncertainty $\sigma$. In the case of energy predictions, this uses the energy error and uncertainty per structure, and in the case of force predictions, it uses the root mean square error and mean uncertainty of all three force components of one atom. $\overline{x}$ denotes the mean over the whole dataset. The value of the correlation coefficient can range from -1 to 1, corresponding to perfect anti-correlation and perfect correlation, respectively, and 0 meaning no correlation at all. The second option, which we are exploring, is the relative log-likelihood (RLL), proposed by Kellner et al. [Kellner2024/10.1088/2632-2153/ad594a]:

### v2

@@ II Methods
The custom-built models consisted of two message-passing layers, with 32 channels and a maximum tensor order of $l=1$. The multi-layer perceptron in the output blocks of the final layer had 16 nodes in the hidden layer of each output block and used the default SiLU gate. The radial cutoff was set to $6.0\text{\,}\AA$ , resulting in an effective field of view of $12.0\text{\,}\AA$ . The isolated atom energies were always set to the ones specified in the respective datasets. The compositional differences between the committee types lead to different sizes of the full training dataset used for each model. To keep the total number of optimization steps consistent between the different types of models, we adapted the training parameters in our setups. This also means that each member of the naive committee received as many optimization steps as the MHC. When investigating different training set sizes, we again ensured that the number of optimization steps remains constant for all training sets. A weighted mean squared error with a force-to-energy weight ratio of 100 to 1 was used for model optimization. For the final 25% of the training, the Stochastic Weight Averaging approach with default settings was used to further optimize the energy predictions. The 3BPA models were trained for 5000 optimization steps, the water models for 2200 steps, and the rMD17 models for 16,000 steps. The output blocks of the adapted foundation model were trained with the same basic parameters for a total of 800,000 steps. To condense the MPtrj dataset, we used an iterative QbC workflow [Schran2020/10.1063/5.0016004]. In each iteration, we sorted the structures based on the maximum disagreement of their force components and added the 100 highest ranking structures to the training set. Afterward, the output heads were retrained for a reduced number of optimization steps.
III Results
In this section, we demonstrate the power of the MHC methodology by applying it to a series of systems and models of increasing complexity. We start with a gas-phase molecule, move to the condensed phase with liquid water, explore chemical space with a model for multiple organic molecules, and finally enhance a foundation model with uncertainty prediction. For the upcoming results, it should be noted that in every calculation, the model’s hyperparameters can influence the results in many unintended, subtle ways, for example the shape of the correlation distribution. We have kept the hyperparameters as consistent as possible between models to minimize these factors.
3BPA
As a first step, we show that the MHC can be used to estimate the uncertainty of MLP predictions. To this end, we test and compare the two different options for subsampling the MHC training set. We then compare these against the baseline of a naive committee, comprising individual full MACE models each trained on randomly chosen subsets of the training data. The aim for the MHC disagreement is to be as close to the naive committee disagreement as possible, despite sharing most of the network’s trainable weights and therefore having less room for divergence between the committee members. For these first tests, we use the 3BPA (3-(benzyloxy)pyridin-2-amine) dataset [Kovacs2021/10.1021/acs.jctc.1c00647], which contains structures of the drug-like 3BPA molecule from MD trajectories at different temperatures. In the 3BPA molecule, shown in the top left panel of Figure 2, the most important degrees of freedom are the torsions along the bonds connecting the two six-membered rings. The training data is based exclusively on $300\text{\,}\mathrm{K}$ structures, whereas the test data originates from sampling at $1200\text{\,}\mathrm{K}$ and therefore reaches a broader region of configuration space than the training set. Figure 2 shows the correlation between the actual error in force or energy per atom and the corresponding standard deviation of the committee predictions. The distribution of the values is shown as a two-dimensional histogram using hexagonal bins on a logarithmic scale. In order to map to the true generalization error, these standard deviations were scaled in postprocessing using a scaling factor calculated from the validation set [Imbalzano2021/10.1063/5.0036522], as detailed in the Methods section. The individual scaling factors $\alpha$ are given in the upper left corner of every panel of Figure 2. To better illustrate the overall trend, we also bin the distribution along the uncertainty axis to contain an equal number of data points in each bin and calculate the mean error for each bin, shown as orange lines. A reference gray line indicates perfect correlation between errors and uncertainty. In an attempt to quantify the correlation, we calculated the Pearson correlation coefficient [Pearson1895/10.1098/rspl.1895.0041] $r$, which can be found in each panel below the scaling factor. We see a significant spread of the distribution surrounding the ideal line, while the binned averages follow the optimal correlation closely. For all models investigated, there are data points with high uncertainty but small error, which is not an issue, as a model can produce an accurate prediction “by accident”, despite high uncertainty. Importantly, there are no instances where the model’s error is high while the uncertainty is low, which would be a clear indication of an unreliable uncertainty estimator. For all of our committee models, there is a clearly empty zone in the bottom right corner, especially for forces, which means that an increased error will always be indicated by an increased uncertainty.
Figure 2: The correlation between the RMSEs of the models and the scaled committee disagreements for forces per atom and energy of the whole molecule. The scaling parameter $\alpha$ and the Pearson correlation coefficient is given in the top left corner of each subplot. The orange line indicates the binned average RMSE for all atoms or structures for forces and energies, respectively, binned along the committee disagreement. All bins contain the same number of data points and are, therefore, not of equal size. The gray line shows perfect correlation between uncertainty measure and actual error. The 3-(benzyloxy)pyridin-2-amine (3BPA) molecule used in this analysis is shown in the inset of the first plot.
Finally, we use the 3BPA system to analyze the prediction accuracy of the different model types. The evolution of the force and energy RMSEs shown in Figure 3 illustrates that the naive committee consistently outperforms the MHCs in prediction accuracy. However, this is expected, as the naive committee has almost 8 times as many trainable parameters as the MHCs due to the full independence of every committee member. Comparing the two different versions of MHCs shows a slight advantage in terms of accuracy for the one with overlapping training sets. In this version, each multi-head training set consists of 80% of the training data for each head for the overlapping committee and just 1/8 of the set for the disjoint committee. The training set for each head comprises 80% of the training data for the overlapping committee and just 12.5% of the training data for the disjoint committee. We conclude that the smaller size of the dataset available to each head in the disjoint committee leads to this small discrepancy. In the bottom row of Figure 3, the difference between the error of the whole committee and the average error of the committee members is shown. For the naive committee, the committee consistently outperforms the individual members due to the increased number of trainable parameters. For the overlapping committee, the differences are small and independent of the number of training structures. The improvement in performance typically associated with committee MLPs remains absent, as the MHC is conceptually similar to adding a dropout layer to the output blocks. While classic dropout layers set the output of random nodes to zero during training, the MHC does so for all nodes connected to certain output heads. For every structure in the training set, the same nodes are consistently removed in every epoch of the training process. It would be unreasonable to expect a significant performance boost from combining multiple of such predictions. Therefore, no significant improvement in predictive power should be expected. Contrary to this expectation, the disjoint model shows a noticeable improvement when introducing the committee. This improvement stems from the number of training examples each output head sees during training. Especially for small training datasets, the subsets for each head are extremely sparse, negatively affecting the prediction accuracy. With a large overall training set size, this sparsity decreases, leading also to more similar single and committee predictions.
Liquid water
Figure 4: Correlation of committee force disagreement and RMSE for each atom in a 64-molecule box of liquid water. The top row shows results for structures with classical nuclei — the same distribution as the training data. The bottom row shows results for structures with quantum nuclei — a distribution different from the training data. In each panel, the Pearson correlation coefficient $r$ is given in the top right corner.
rMD17
While each of the previous test cases focused on a single molecular system, in this section, we expand our investigation to cover a more diverse set of molecules within one model to probe uncertainty predictions across chemical compound space. To test this, we employ the rMD17 [Christensen2020/10.1088/2632-2153/abba6f] dataset comprising MD structures sampled at $500\text{\,}\mathrm{K}$ of 10 different organic molecules (consisting of H, C, N, and O atoms), which are shown in Figure 5. We randomly selected 50 structures from each of the 10 molecules to form our training set and 1000 structures of each molecule for separate per-species test sets. The unscaled mean and standard deviation of the uncertainties are shown in Figure 5 against the force errors for the multi-head and naive committees for all 10 molecules individually, to show how the model can handle predictions of different complexities. For an easier comparison, we plot the binned averages of all three approaches in one subplot. The shaded area indicates the standard deviation of the data in each bin to illustrate the spread of the data. The magnitudes of the uncertainty are dependent on the molecule under investigation, but the general trends are very consistent, as are the comparisons between the committee types. Analogously to our previous findings, the scales of the three curves differ substantially, while the correlations remain fairly similar for all three committee types. As expected, the naive committee shows both the highest level of uncertainty and the highest Pearson correlation coefficient due to the large differences between its members, and the overlapping MHC shows the lowest, as its members are the least diverse. We also conducted a correlation analysis using the relative log-likelihood method [Kellner2024/10.1088/2632-2153/ad594a], coming to the same conclusions (details in the SI, Section S2. Unfortunately, the correlation between energy uncertainty and energy errors is substantially weaker. The Pearson correlation coefficient calculated over the full test set is only 0.37 for both the naive committee and the disjoint MHC and 0.29 for the overlapping MHC. We discuss the energy uncertainty in further detail in the SI, Section S1. We also use the rMD17 dataset to examine committee disagreement for unknown atomic systems, as this will be relevant in the context of foundation models. Therefore, we remove one molecule from the training data and train the committees on the remaining 9 molecules. When we examine the committee disagreement on a test set of the molecule that was taken out of the training data, we find that the uncertainty still correlates well with the errors but can be distinctively low for certain atoms. We attribute this to the local environment of these atoms, which is similar to the local environment of atoms present in the training data. We discuss these results in more detail in the SI S3. Importantly, we find that our previous conclusions are also valid for heterogeneous datasets and that committee disagreement works well with large, diverse training sets and out-of-domain test structures.
Figure 5: Correlation between the unscaled committee disagreement and the force RMSE for individual atoms in the systems. The curves are plotted individually for all 10 different molecules in the rMD17 dataset and the structure of the respective molecule is shown above the plots. The curves show the average force RMSE binned along the committee disagreement, as the orange curves in Figure 2 and 4. The shaded areas around the lines are the standard deviations within each bin. The Pearson correlation coefficient $r$ is given above the correlation plots in the color of the respective model.
MACE MP-0 foundation model
After showing that the MHC provides a reliable measure of prediction uncertainty, we adapt it to a foundation model. We start from the MACE-MP-0b foundation model [Batatia2023/10.48550/arXiv.2401.00096] trained on the MPtrj dataset [Deng2023/10.1038/s42256-023-00716-3], which contains 146k crystalline structures calculated using density functional theory at the PBE+U [Perdew1996/10.1103/PhysRevLett.77.3865] level. We then equip this foundation model with an MHC by adding eight new output heads with a random initialization of weights to the existing pre-trained head. As with custom-build models, the committee prediction is the mean of all new output heads, excluding the original head. When training the MHC output heads, we leave the weights of the AED block and the original head fixed. To obtain a more compact training set for the MHC, we use an iterative QbC workflow based on its committee disagreement [Schran2020/10.1063/5.0016004]. This reduces the original dataset to 8,000 structures — just 5% of its original size. This condensed training dataset is divided across the eight committee heads using the “disjoint” distribution, resulting in 1,000 training structures for each head. By using a much smaller training set and taking advantage of the pre-trained model, the computational cost of training is greatly reduced. Overall, we are adding roughly 15,000 parameters to the model, accounting for less than 0.2% of the total model size. It took 37 hours on an NVIDIA Hopper H100 GPU to execute the 800,000 optimization steps of the whole MHC, a small fraction of the 2,600 GPU hours of the original model (trained on NVIDIA A100 GPUs across multiple nodes) [Batatia2023/10.48550/arXiv.2401.00096]. Especially for the smaller foundation models, this training effort can also be performed on consumer-grade GPUs.
Figure 6: Correlation between the committee disagreement and the force error for the modified MACE MP-0b model, where the output heads were trained on a reduced sample of the MPtrj dataset containing 8000 structures. The first subplot is on test data from the MPtrj dataset, which is not included in the reduced dataset. For the two subsequent plots, 10,000 structures from the NVT and NpT dataset of the OMAT database were used. As in the previous figures, the orange lines indicate the average force RMSE, binned along the direction of the uncertainty measure. The Pearson correlation coefficient $r$ is given in the top left corner of each panel.
For evaluation, we used 10,000 out-of-distribution structures taken from the NVT and NpT OMAT datasets [Barroso-Luque/10.48550/arXiv.2410.12771], which are calculated with the same reference method as the MPtrj dataset, and 10,000 structures from the MPtrj dataset not selected during QbC. For the OMAT datasets, a very small number of predictions had an extremely high error for all models, including the original MACE-MP-0 model. Due to the nature of the Pearson correlation coefficient and the root mean square errors, these few predictions dominate the final results. Therefore, we decided to exclude all force components for which the error of the original MACE-MP-0 model is higher than $5\text{\,}\mathrm{e}\mathrm{V}\mathrm{/}\AA$ and all energy predictions with an energy error higher than $1\text{\,}\mathrm{e}\mathrm{V}$. This amounts to a total of 18 energy values and 764 force components, or 0.019% of the combined OMAT test sets. Figure 6 shows the correlation between the actual force errors per atom and the uncertainty prediction for the three datasets in question. Compared to previous cases, the results are more spread out, as both the training and test data are much more diverse. In the case of OMAT, this means that many of the systems in the test set are not present in the training data at all. The MPtrj dataset, on the other hand, functions as an indicator for the in-data regime, as it is not independent of the training data of the model. The AEDs of the model were trained on the full MPtrj dataset, including the structures of this test set. Overall, the committee disagreement correlates well with the error of the committee prediction and therefore functions as a reliable uncertainty measure even for foundation models. However, unlike with custom-made models, we observe occasional instances where a low uncertainty coincides with a high error. This is most likely due to structures in the test set, which have similar characteristics to some structures in the training set, resulting in uniform predictions by the committee. This also explains why the Pearson correlation coefficient is moderately lower than for custom-made models. In the MPtrj dataset, where the structures of the test set were used to train the original foundation model, but not the committee output heads, this effect is particularly pronounced. However, overall, these instances are isolated enough not to deteriorate the reliability of the MHC uncertainty prediction for foundation models.
An important question that remains is whether the new committee model predictions are still accurate compared to the original, despite being trained on a small fraction of the original dataset. To test whether QbC is advantageous compared to other, computationally cheaper options, we created two alternative training sets. During the QbC run to select the reduced training dataset, it is noticeable that structures with high force components are selected at a much higher rate than structures close to the equilibrium. Therefore, we created a training set of the 8,000 structures with the highest force per atom. Additionally, we also randomly selected 8,000 structures from the initial dataset to form a third training set. Both datasets were used to train MHC models in the same way as the QbC dataset. Figure 7 shows the distribution of the force errors for the original MACE-MP-0 model and the new MHC models. As expected, the original MACE-MP-0 model has the highest prediction accuracy of all models, but the advantage over the re-trained models is small. Among the models trained on less data, the model with the QbC selected data performs the best, but the differences compared to other models are modest. For the MPtrj dataset, the original MACE-MP-0 model has a much stronger advantage over the retrained models than for the OMAT data. This is because the OMAT data is entirely unseen by all models, whereas the MPtrj test set was part of the MACE-MP-0 training dataset, and therefore, this strong performance should be expected. This also explains why the randomly selected subset outperforms the QbC selected set, as we sampled it from the structures not picked during QbC. Therefore, some structures of the randomly selected training set are included in this test set. As we show in the SI S4, condensing the training set even further will eventually lead to a measurable decrease in prediction accuracy.
Overall, we conclude that our approach of adding an MHC to a foundation model preserves its predictive power. This is, on one hand, achieved by leaving the original foundation model — including its output head — intact. On the other hand, we have shown that the added output heads, trained on a condensed training dataset, also largely retain the foundation model’s performance, as most of the predictive capacity lies within its AEDs, which were extensively trained on the full MPtrj dataset. Optimizing the weights of the AEDs further on a condensed training set would degrade the model’s performance.
IV Conclusions
The second part of this work used the MHC in active learning to condense a reduced training set out of the original MPtrj dataset, a widely used training set for foundation models. We selected 8000 out of 146k structures from the MPtrj dataset and used them to train the output heads of an adapted medium-sized MACE-MP-Ob foundation model. The AEDs of the model were left untouched to preserve its expressive capabilities. We showed that the MHC based on foundation models displays a good correlation between committee disagreement and actual force error, even though instances where the uncertainty is underestimated are slightly more common. When comparing the predictions of the original MACE-MP-0b model with the new model, we saw only a small drop in performance, although the output heads were trained on only 5.5% of the full training dataset. We also examined the criteria for selecting the reduced training sets and found that the QbC-selected model performs the best, even though the advantage is moderate. Given the compact nature of the condensed data sets, this opens up the possibility to obtain foundation models trained on expensive high-level electronic structure methods by recalculating the previously condensed training set and optimizing only the output blocks. QbC can help ensure that the performance stays as close as possible to the original foundation model. Overall, the uncertainty estimation of the MHC architecture introduced here further increases the robustness of foundation models.
Acknowledgements.
The authors thank Ilyes Batatia and Gábor Csányi for valuable discussions. H.B. and P.S. acknowledge support from the Charles University Grant Agency, project number 248923, and the International Max Planck Research School for Quantum Dynamics and Control. L.L.S. would like to acknowledge support from the UKRI Critical Mass grant, project reference EP/V062654/1, the Isaac Newton Trust, award number G122390 and Wolfson College, Cambridge. O.M. acknowledges support from the Czech Science Foundation, project No. 21-27987S. C.S. acknowledges financial support from the Royal Society, grant number RGS/R2/242614.
Conflict of Interest Statment
The authors have no conflicts to disclose.
Author contributions
@@ Acknowledgements.
The data that support the findings of this study are openly available in Zenodo at http://doi.org/[doi] (doi will be added when published)
@@
When investigating how well committee disagreement functions as an uncertainty estimation for the true error in the main article, we focused on a graphical evidence and plotted the uncertainty estimate against the actual error. However, there are alternatives to assess uncertainty in a more quantitative way, which we show here. The first is the Pearson correlation coefficient [Pearson1895/10.1098/rspl.1895.0041], which we explained in the Section II of the main paper. For many results in the main paper, the Pearson correlation coefficient was given as an annotation in the figures. The second option, which we are exploring, is the relative log-likelihood (RLL), proposed by Kellner et al. [Kellner2024/10.1088/2632-2153/ad594a]: